OpenStudy (anonymous):

An angle measure of 82 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary.

OpenStudy (anonymous):

divide by \(180\) then stick a \(\pi\) next to it

OpenStudy (anonymous):

Thank you! Is that all?

OpenStudy (anonymous):

I got a really weird number

OpenStudy (anonymous):

you just write \[\frac{82\pi}{180}\] if you like or reduce the fraction to \[\frac{41\pi}{90}\] the decimal will be long

OpenStudy (anonymous):

well ya but it said to round the decimal to the nearest hundreth

OpenStudy (anonymous):

so would it be 1.4?